OpenStudy (anonymous):

Perpendicular to the graph of y= 1/3x+7

OpenStudy (anonymous):

do you know how to find the perpendicular slope?

OpenStudy (anonymous):

I know it is the reciprocal and the opposite of 1/3x right?

OpenStudy (anonymous):

yep so what is the slope?

OpenStudy (anonymous):

-3/1x or -3x

OpenStudy (anonymous):

yep now if you sub that into y=mx +b for m you have a equation for the perpendicular line :)

OpenStudy (anonymous):

I am taking online college and as my homework I put y=-3/1x+7 and my teacher said its wrong, something about the y-intercept??

OpenStudy (anonymous):

i would not know why thats wrong then is there more information to the question?

OpenStudy (anonymous):

hold on ill get it

OpenStudy (anonymous):

The question is:Wright the slope-intercept equation of the line whose graph is described. Perpendicular to the graph of y=1/3x+7 with the y-intercept of (0,5) maybe I have to sub 5 where the 7 is????

OpenStudy (anonymous):

yep :) because 7 is the y-intercept and the question tells us to have a 5 y-intercept :)

OpenStudy (anonymous):

Okay Thanks. I feel dumb now lol.

OpenStudy (anonymous):

its all good i misread all the time is a common mistake it gets the best people sometimes :)

OpenStudy (anonymous):

thanks again!!

OpenStudy (anonymous):

no problem :)